# Simulation of the process $e^{+}e^{-}\rightarrow W^{+}W^{-}$ with the heavy right-handed neutrino exchange at 1 TeV future lepton colliders

### v1

@@ I INTRODUCTION
The process $l^{-}l^{-}\to W^{-}W^{-}$ ($l=e,\mu$) has been discussed in several papers [12, 13], assuming opportunity for future same-sign electron or muon colliders. This process with $t$-channel HNL exchange is possible, if heavy neutrinos have a Majorana nature. In this process the lepton number violates by two units, therefore it has no backgrounds in the Standard Model. Using our experimental procedures we can well reproduce the results obtained in these papers. However, the same-sign lepton colliders are not considered as the possible option due to technical reasons. Therefore we study the similar process for the case of the opposite-sign beams.
In contrast to processes with direct HNL production [14, 15], the $e^{+}e^{-}\rightarrow W^{+}W^{-}$ process is sensitive to contributions of HNLs with very large masses, $M(N)\gg\sqrt{s}$. Within the framework of the Seesaw type-I model, HNL masses can be large, therefore it is important to have experimental accessibility to the large HNL mass region. Since heavy neutrinos are right-handed and active neutrinos are left-handed, interference of the SM and HNL diagrams results in specific angular distributions in the $W$ boson production and decays. We use these distributions within the extended likelihood method to evaluate the HNL contribution to the process and to obtain upper limits on the HNL mixing parameters.
Current experimental upper limits on the mixing parameters $|V_{lN}|$, $l=e,\mu,\tau$ have been discussed in [16]. There are strict upper limits $|V_{eN}|^{2}\lesssim 10^{-5}$ for the HNL masses less than the $Z$ boson mass. However in the region $M(N)>100~{}\text{GeV}$, the upper limits, mostly obtained by the LHC, are weak, $|V_{eN}|^{2}\gtrsim 10^{-2}$ for $M_{N}<1~{}\text{TeV}$ and $|V_{eN}|^{2}\gtrsim 1$ for $M_{N}>1~{}\text{TeV}$. The latest results of searches for HNL performed by the LHC collaborations are presented in [17, 18, 19, 20, 21]. Future lepton colliders can significantly improve the upper limits in the large HNL mass region up to ${\cal{O}}(10^{4}-10^{5})~{}\text{GeV}$.
@@ II.4 Angular distributions
The $e^{+}e^{-}\rightarrow W^{+}W^{-}$ process is described by five angles $\theta,\theta_{1},\theta_{2},\phi_{1},\phi_{2}$ as shown in Fig. 5. Three out of them, $\theta,\theta_{1},\theta_{2}$ are used to distinguish the SM+HNL contribution from the SM-only contribution. The production angle $\theta$ is the angle between the $W$ boson direction and the $e^{-}$ beam direction in the laboratory system. The decay angles $\theta_{1}$ and $\theta_{2}$ are the angles between the jet direction and the $e^{-}$ beam direction in the $W$ boson rest frame. $\theta_{1}$ is the decay angle of $W$ boson with the larger angle with the $e^{-}$ beam direction and $\theta_{2}$ is the decay angle of the second $W$ boson. Because of the right-handed polarization of heavy neutrino, the distributions of the variables $\theta,\theta_{1},\theta_{2}$ for the SM and HNL contributions are different. Therefore we can use these distributions to estimate the significance of the HNL contribution. The distributions of angles $\phi_{1}$ and $\phi_{2}$ between the production and decay planes for the signal and background events are very similar and are not used in this study.

### v2

@@ I INTRODUCTION
The process $l^{-}l^{-}\to W^{-}W^{-}$ ($l=e,\mu$) has been discussed in several papers [12, 13, 14], assuming opportunity for future same-sign electron or muon colliders. This process with $t$-channel HNL exchange is possible, if heavy neutrinos have a Majorana nature. In this process the lepton number violates by two units, therefore it has no backgrounds in the Standard Model. Using our experimental procedures we can well reproduce the results obtained in these papers. However, the same-sign lepton colliders are not considered as the possible option due to technical reasons. Therefore we study the similar process for the case of the opposite-sign beams.
In contrast to processes with direct HNL production [14, 15, 16], the $e^{+}e^{-}\rightarrow W^{+}W^{-}$ process is sensitive to contributions of HNLs with very large masses, $M(N)\gg\sqrt{s}$. Within the framework of the Seesaw type-I model, HNL masses can be large, therefore it is important to have experimental accessibility to the large HNL mass region. Since heavy right-handed neutrinos have large mass, interference of the SM and HNL diagrams results in specific angular distributions in the $W$ boson production and decays. We use these distributions within the extended likelihood method to evaluate the HNL contribution to the process and to obtain upper limits on the HNL mixing parameters. At the region $\sqrt{s}\sim 200-500~{}\text{GeV}$ the SM $e^{+}e^{-}\to W^{+}W^{-}$ cross section is $\sim 1-2~{}\text{pb}$ and the $t$-channel HNL exchange contributions are too small for experimental searches. In the paper [17] virtual HNL loop contributions were calculated and found to be also very small in this $\sqrt{s}$ region.
Current experimental upper limits on the mixing parameters $|V_{lN}|$, $l=e,\mu,\tau$ have been discussed in [18, 19]. There are strict upper limits $|V_{eN}|^{2}\lesssim 10^{-5}$ for the HNL masses less than the $Z$ boson mass. However in the region $M(N)>100~{}\text{GeV}$, the upper limits, mostly obtained by the LHC, are weak, $|V_{eN}|^{2}\gtrsim 10^{-2}$ for $M_{N}<1~{}\text{TeV}$ and $|V_{eN}|^{2}\gtrsim 1$ for $M_{N}>1~{}\text{TeV}$. The latest results of searches for HNL performed by the LHC collaborations are presented in [20, 21, 22, 23, 24]. Future lepton colliders can significantly improve the upper limits in the large HNL mass region up to ${\cal{O}}(10^{4}-10^{5})~{}\text{GeV}$.
@@ II.4 Angular distributions
The $e^{+}e^{-}\rightarrow W^{+}W^{-}$ process is described by five angles $\theta,\theta_{1},\theta_{2},\phi_{1},\phi_{2}$ as shown in Fig. 5. Three out of them, $\theta,\theta_{1},\theta_{2}$ are used to distinguish the SM+HNL contribution from the SM-only contribution. The production angle $\theta$ is the angle between the $W$ boson direction and the $e^{-}$ beam direction in the laboratory system. The decay angles $\theta_{1}$ and $\theta_{2}$ are the angles between the jet direction and the $e^{-}$ beam direction in the $W$ boson rest frame. $\theta_{1}$ is the decay angle of $W$ boson with the larger angle with the $e^{-}$ beam direction and $\theta_{2}$ is the decay angle of the second $W$ boson. Because of the large mass of heavy neutrino, the distributions of the variables $\theta,\theta_{1},\theta_{2}$ for the SM and HNL contributions are different. Therefore we can use these distributions to estimate the significance of the HNL contribution. The distributions of angles $\phi_{1}$ and $\phi_{2}$ between the production and decay planes for the signal and background events are very similar and are not used in this study.